OpenStudy (anonymous):

Some farmers use a circular irrigation method. An irrigation arm acts as the radius of an irrigation circle. How much more land is covered with an irrigation arm of 300 ft than by an irrigation arm of 250 ft?

OpenStudy (isaidavila):

I think you should rest the áreas of a circle of 300 ft radius and other of 250ft radius A1=π(300)^2 A2=π(250)^2